# Modeling error correction with Lindblad dynamics and approximate channels

### v2

@@ V.5 Purely ZZ crosstalk
$h=0,\qquad T_{1}=\infty,\qquad T_{2}=\infty.$ (42)
@@ V.6 Lower crosstalk noise
For these parameters, as can be seen in Fig. 3, the regular decoder is advantageous, and so we will only use the regular decoder throughout this section. In Fig. 6 we present $\eta$ for the different approximations, on the average of the initial states in Eq. (9), and note that the individual states behave similarly. We can see that the composite-channel approximation, as well as the Pauli approximation, are very good approximations within the short-time limit of the dynamics, though they begin to lose accuracy at $t\sim 6\mu$s.
@@ VI Summary and Outlook
This work is focused on the importance of physically-motivated noise dynamics when determining the success rate of an error correction code to recover logical states subject to noisy dynamics. The Lindblad model studied in this work includes the dominant idle errors in superconducting quantum computing systems [20], or can be modified in order to better model those  [^5] We note that the effect of frequency shifts caused by stochastic charge-parity fluctuation in superconducting qubits [20], can be absorbed into each qubit’s $h_{i}$ value and averaged over the possible frequencies., and could be relevant for other systems as well [64, 65, 66, 67, 68, 69]. As mentioned in Sec. II, in some quantum devices the qubit parameters vary from qubit to qubit. The natural question that arises is how an inhomogeneity in the dynamical parameters would affect the conclusions of our study. In App. B we have considered the five-qubit code dynamics with some spread in the parameters (relevant to superconducting qubits), and we see that the error-correction dynamics of the spatially inhomogeneous system could still be qualitatively similar to the dynamics observed with uniform qubits.
Overall, we have determined that for a Lindbladian of the form presented in Sec. II, the composite-channel approximation (constructed explicitly to account for 2Q noise) is a very good approximation for relevant noise parameters and the time scales that we study. We show that the composite-channel approximation holds up to time scales for which the non-commutativity of the individual channels becomes nonnegligible. This approximation will also hold for other systems with similar ratios between noise parameters, by scaling the duration of the noise channel accordingly. Modeling other coupling terms can be more subtle due to noncommuting terms possibly contributing in shorter time scales, as in the case of an XY coupling term, which we briefly consider in App. H. This understanding would be an important a guideline when modelling more general setups than the one considered here, e.g., when incorporating the dynamics during gates involving parity-check qubits. We have also considered the effect of choosing a reference frame, such as the non-symmetric reference frame often used with superconducting qubits hardware, the different ways the Pauli approximation can be performed in this case, and the affect it has on the performance of the approximation.
@@
Therefore, a fundamental open question is how our observations would change with circuit-based non-Pauli dynamics of error correction, and what would be the relevant effects with larger codes with an increasing distance, and some randomness in the parameters. In particular, how the error correction threshold and logical error rates scale and change when treated within the different approximations. This direction could be nontrivial to pursue since it would require the simulation of the error correction procedure with large codes. This may be difficult with a dynamical simulation, but perhaps feasible with the composite-channel approximation, and in particular, using tensor-network methods that are suitable for the relevant states [20]. An understanding of this question may be useful to shed a light on the feasibility of error correction under realistic noise and could possibly contribute to its realization.
@@ C.3 Dynamical noise on a physical qubit
We compared the same noise channel on the physical qubit and on the logical qubit. Remembering that the physical qubit is much simpler than the logical qubit, we consider also the case where simulating the dynamical simulation of the noise is possible for the physical qubit, and not for the logical qubit. So when comparing to the curve of the dynamical simulation of the noise on the physical qubit, the pseudo-thresholds of the composite approximations do not change, but the Pauli noise pseudo-threshold becomes $5.84\,\mu$s. This means that only a small part of the mismatch was caused by the inaccuracy of the Pauli approximation on the physical qubit.
@@ Appendix G Pauli approximation methods
$U_{3}=e^{-\frac{1}{2}i\left(\sigma^{x}\sigma^{x}+\sigma^{y}\sigma^{y}\right)Jt}=\left(\begin{matrix}1&0&0&0\\ 0&\cos Jt&-i\sin Jt&0\\ 0&-i\sin Jt&\cos Jt&0\\ 0&0&0&1\end{matrix}\right).$ (70)
@@
Now, while the commutator in this case is of second order in the noise parameters as in the case of $\mathcal{U}^{zz}_{\text{2Q}}$ with the 1Q damping, for the studied range of noise parameters, the $\mathcal{U}^{xy}_{\text{2Q}}$ commutators can be significantly larger since $J\gg 1/T_{1}$. Detailed explanations can be found in App. I. And so we would expect a more significant error of the composite-channel approximation in this case. And indeed we can see in Fig. 13 that for $t\gtrsim 2\,\mu$s the composite-channel approximation no longer holds, not even qualitatively.
@@ I.2 Failure for ZZ crosstalk and large damping noise
The composite-channel approximation fails to properly incorporate the non-trivial commutation relations between the noise terms. In Fig. 13 we see that the composite relation fails at $t\simeq 2.5\mu s$ in the presence of $\mathcal{U}_{\text{2Q}}^{xy}$ noise. In this case, $J=2\pi\times 30$ kHz and so the commutator is proportional to $(Jt)^{2}\simeq 0.2$, and is indeed nearing one. On the other hand, in Fig. 2 where the non commuting terms are $\mathcal{U}_{2Q}^{z}$ and the damping noise, the composite-channel approximation holds very well. In this case, the noise parameters, and specifically the damping parameters are relatively small. For $\zeta=-2\pi\times 30\,$kHz and $T_{1}=150\mu s$, the commutator for $t\simeq 2.5\mu s$ is proportional to $\zeta t\frac{t}{T_{1}}\simeq 0.0067$. When analyzing the non realistic case, where $T_{1}=5\mu s$ giving rise to $\zeta t\frac{t}{T_{1}}\simeq 0.2$, then as can be seen in Fig. 14, the composite-channel approximation differs from the dynamical simulation more significantly than for smaller $\zeta$ as in Fig. 2, but the deviation is not qualitative as in Fig. 13.

### v3

@@ V.5 Purely ZZ crosstalk
$h=0,\qquad T_{1}=T_{\phi}=T_{2}=\infty.$ (42)
@@ V.6 Lower crosstalk noise
For these parameters, as can be seen in Fig. 3, the regular decoder is advantageous, and so we will only use the regular decoder throughout this section. In Fig. 6 we present $\eta$ for the different approximations, on the average of the initial states in Eq. (9), and note that the individual states behave similarly. We can see that the composite-channel approximation, as well as the Pauli approximation, are very good approximations within the short-time limit of the dynamics, though they begin to lose accuracy at $t\sim 6\mu$s, consistent with expected (small but not negligible) contributions from coherent accumulation of phases (1Q and 2Q) and noncommuting terms.
@@ VI Summary and Outlook
This work is focused on the importance of physically-motivated noise dynamics when determining the success rate of an error correction code to recover logical states subject to noisy dynamics. The Lindblad model studied in this work includes the dominant idle errors in superconducting quantum computing systems [22], or can be modified in order to better model those  [^5] We note that the effect of frequency shifts caused by stochastic charge-parity fluctuation in superconducting qubits [22], can be absorbed into each qubit’s $h_{i}$ value and averaged over the possible frequencies.. The Markovian decoherence and qubit crosstalk that we focused on here constitute in many physical systems the leading-order noise terms [68, 69, 70, 71, 72, 73]. As noise is reduced on qubits and control techniques improve, non-Markovian modeling of the environment may become important [74, 75, 76, 77, 78, 79] and it would be interesting to see how approximations of error-correction dynamics perform.
As mentioned in Sec. II, in some quantum devices the qubit parameters vary from qubit to qubit. The natural question that arises is how an inhomogeneity in the dynamical parameters would affect the conclusions of our study. In App. B we have considered the five-qubit code dynamics with some spread in the parameters (relevant to superconducting qubits), and we see that the error-correction dynamics of the spatially inhomogeneous system could still be qualitatively similar to the dynamics observed with uniform qubits.
Overall, we have determined that for a Lindbladian of the form presented in Sec. II, the composite-channel approximation (constructed explicitly to account for 2Q noise) is a very good approximation for relevant noise parameters and the time scales that we study. We show that the composite-channel approximation holds up to time scales for which the non-commutativity of the individual channels becomes nonnegligible. This approximation will also hold for other systems with similar ratios between noise parameters, by scaling the duration of the noise channel accordingly. Modeling other coupling terms can be more subtle due to noncommuting terms possibly contributing in shorter time scales, as in the case of an XY coupling term, which we briefly consider in App. H. This understanding would be an important guideline when modelling more general setups than the one considered here, e.g., when incorporating the dynamics during gates involving parity-check qubits. We have also considered the effect of choosing a reference frame, such as the non-symmetric reference frame often used with superconducting qubits hardware, the different ways the Pauli approximation can be performed in this case, and the affect it has on the performance of the approximation.
@@
Therefore, a fundamental open question is how our observations would change with circuit-based non-Pauli dynamics of error correction, and what would be the relevant effects with larger codes with an increasing distance, and some randomness in the parameters. It will be interesting to identify which properties are common to all codes, and which properties are unique to different codes. In particular, how the error correction threshold and logical error rates scale and change when treated within the different approximations, and whether the possible accumulation of coherent logical errors across syndrome cycles observed in Sec. V.8 remains relevant with codes of higher distance. This direction could be nontrivial to pursue since it would require the simulation of the error correction procedure with large codes. This may be difficult with a dynamical simulation, but perhaps feasible with the composite-channel approximation, and in particular, using tensor-network methods that are suitable for the relevant states [22]. An understanding of this question may be useful to shed a light on the feasibility of error correction under realistic noise and could possibly contribute to its realization.
@@ C.3 Dynamical noise on a physical qubit
In all the above, we compared the same noise channel on the physical qubit and on the logical qubit. Remembering that the physical qubit is much simpler than the logical qubit, we consider also the case where simulating the dynamical simulation of the noise is possible for the physical qubit, and not for the logical qubit. In such a case it will be beneficial to identify how much of the discrepancy in the pseudo-threshold is caused by the approximation of the noise on the logical qubit, and how much is caused by the approximation of the noise on the physical qubit, since the latter can be eliminated by exact simulation of the noise. When we simulate dynamical noise on the physical qubit, and approximated noise on the logical one, the pseudo-thresholds of the composite approximations do not change, but the Pauli noise pseudo-threshold becomes $5.84\,\mu$s. This means that only a small part of the mismatch was caused by the inaccuracy of the Pauli approximation on the physical qubit.
@@ Appendix G Pauli approximation methods
$U_{3}=e^{-i\left(V^{xy}_{12}\right)t}=\left(\begin{matrix}1&0&0&0\\ 0&\cos Jt&-i\sin Jt&0\\ 0&-i\sin Jt&\cos Jt&0\\ 0&0&0&1\end{matrix}\right).$ (70)
@@
Now, while the commutator in this case is of second order in the noise parameters as in the case of $\mathcal{U}^{zz}_{\text{2Q}}$ with the 1Q damping, for the studied range of noise parameters, the $\mathcal{U}^{xy}_{\text{2Q}}$ commutators can be significantly larger since $J\gg 1/T_{1}$ (a more detailed analysis can be found in App. I), and so we would expect a more significant error of the composite-channel approximation in this case. Indeed, we can see in Fig. 13 that for $t\gtrsim 2.5\,\mu$s the composite-channel approximation no longer holds, not even qualitatively.
@@ I.2 Failure for ZZ crosstalk and large damping noise
The composite-channel approximation fails to properly incorporate the non-trivial commutation relations between the noise terms. In Fig. 13 we see that the composite relation fails at $t\simeq 2.5\mu s$ in the presence of $\mathcal{U}_{\text{2Q}}^{xy}$ noise. In this case, $J=2\pi\times 30$ kHz and so the commutator is proportional to $(Jt)^{2}\simeq 0.22$, and is indeed nearing one. On the other hand, in Fig. 2 where the non commuting terms are $\mathcal{U}_{2Q}^{z}$ and the damping noise, the composite-channel approximation holds very well. In this case, the noise parameters, and specifically the damping parameters are relatively small. For $\zeta=-2\pi\times 30\,$kHz and $T_{1}=150\mu s$, the commutator for $t=2.5\mu s$ is proportional to $|\zeta|t\frac{t}{T_{1}}\approx 0.008$. When analyzing the non realistic case, where $T_{1}=5\mu s$ giving rise to $|\zeta|t\frac{t}{T_{1}}\approx 0.24$, then as can be seen in Fig. 14, the composite-channel approximation differs from the dynamical simulation more significantly than for smaller $|\zeta|$ as in Fig. 2, but the deviation is not qualitative as in Fig. 13.